### v1

# Chain-structured neural architecture search for financial time series forecasting

Abstract
We compare three popular neural architecture search strategies on chain-structured search spaces: Bayesian optimization, the hyperband method, and reinforcement learning in the context of financial time series forecasting.
1 Introduction
Deep neural networks have been very successful in a wide variety of tasks over the last two decades. In large part their success is attributed to their ability to perform very well without major manual feature engineering required when compared to more classical techniques. The hierarchical structure of the networks weighs and extracts important features automatically from the data during the learning process [GBC16]. However, the exact architecture of the neural network still has to be prescribed manually by the user. This led to the development of so-called auto-ML techniques that aim to automate this process. In the context of deep neural networks, auto-ML has a very large overlap with neural architecture search (NAS), itself having a large overlap with hyperparameter optimization.
A lot of research has been done in NAS in recent years, see [${}^{+}$23] for an overview and insights from over 1000 papers. However, most work focused on computer vision or natural language applications, with less investigation into architectures for analyzing time series data. In this work, we attempt to bridge this gap by evaluating and comparing the performance of three popular simple NAS strategies on 3 distinct yet similar time series datasets.
A seminal survey article on NAS [EMH19] breaks down the search into 3 components: search spaces, search strategies and performance estimation strategies. Once a search space is set up, an algorithm following the search strategy explores the search space looking for the best neural network configuration, evaluated and eventually selected by the performance estimation strategy. In Section 2, we discuss the exact practical problem from project partner Predictive Layer SA that our neural networks were trying to solve. We discuss the search spaces we set up for NAS in Section 3, the search strategies in Section 5 and the performance estimation together with the general challenges of our problem in Section 4. Our results are summarized and discussed in Section 6 and Section 7 respectively. Finally, we outline possible future directions for this research in Section 8.
2 Data and problem formulation
This project was a collaboration with Predictive Layer SA, who provided real-world customer data and prediction requirement. The data under consideration are financial multivariate daily time series in tabular form, each row corresponding to one day and each column a separate feature. The task consists of predicting whether the target feature will increase or decrease 5 days (or 10 days, depending on the dataset) in the future. Thus the problem is essentially binary classification. The models output a number between 00 and $1$ as output, interpreted as probability that the target feature will grow on a given day.
The challenge comes with financial markets being notoriously hard to predict [HH19]. Our three source datasets that we report on in Section 6 are for Japanese, German, and US bonds and come with an additional challenge of very high number of features (around 1000 each), comparable with the number of training examples (under 4000), which is very low in a deep learning setting [GBC16], especially relatively to the number of features. To circumvent the high dimensionality, the number of features was reduced approximately by a factor of $3$ by removing features in the prepared datasets ‘time-derived’ from original features, such as by lagging original features by fixed time steps or taking the mean over the same time steps. Such time-derived features are standard in the industry [HH19], but are fundamentally used because the algorithms applied to the data can only take in one row of data as input. All the more sophisticated neural network architectures take in a sequence of rows of data, making time-derived features redundant. See Section 3 for more details. Additionally, principal component analysis (PCA) was applied to further reduce the number of features. Based on the scree plot coming from PCA analysis, two breakpoints were tested: the point of highest curvature (the usual choice) and the point where the graph flattens, retaining most of the information. Thus for each of the 3 source datasets, we obtained 4 variations: no modification, time-derived features removed and 2 PCA transformations, resulting in 12 datasets total.
3 Architecture types and their search spaces
In NAS, the search spaces and search strategies go hand in hand and are designed together essentially at the same time. The simplest among the search spaces are chain-structured. Search spaces with a chain structure feature a straightforward architectural topology: a sequential chain of operational layers. These configurations frequently utilize cutting-edge manually designed architectures as their foundational framework, trying to find the best configuration for the specific dataset at hand by essentially hyperparameter optimization. Despite their simplicity, chain-structured search spaces often yield very good results [${}^{+}$23]. The biggest downsides of this approach when compared to the more complicated cell-based and one-shot approaches are its limited flexibility and scalability when applied to very large datasets. While being limited to a pre-defined overall structure prevents discovering truly novel architectures, in practice well-designed and well-understood architectures usually perform best [${}^{+}$23]. Scalability becomes important when dealing with large image and natural language datasets that most research on NAS focuses on, however in the case of time series and especially our small datasets, our neural networks train in seconds. Therefore scalability was not an issue for us, and thus we focused on chain-structured search spaces.
For the foundational frameworks the search spaces were built upon, we concentrated on simple well-known architecture types, like feedforward (FFNN, also known as multi-layer perceptron), convolutional (CNN), and recurrent (RNN) networks. This was done both for their proven perforamnce in time-series context and due to them having fewer parameters than most cutting-edge architectures, which is more appropriate when dealing with very little data as in our case. However, we also examined the state of the art Temporal Fusion Transformer (TFT) architecture for comparison (see below). Each of these architectures was adapted to the problem at hand and for each of them a search space of hyperparameters was set up to determine the exact configuration. The search strategies discussed in Section 5 then optimized each of these separately, and the best-performing ones were compared. This approach is similar to what [ZL17] did with a single type of CNN for image classification.
3.1 Feedforward networks
Feedforward networks have the advantage of being the simplest type of deep neural networks. The main disadvantage for our problem comes from the fact that they cannot take the time-dimension into account explicitly. Feedforward networks take a row of numbers as input, one row of our datasets corresponding to one day information for various time series. Implicitly, the time information is taken into account in the form of delayed (or averaged) features – a column in the dataset that is essentially the same as another column with an added delay of a couple days (or averaged over a couple days, respectively), a standard practice in the field [HH19].
As far as NAS is concerned, our hyperparameters for optimization are the number of hidden layers and the number of units per hidden layer. Dropout layers were used for regularization and the dropout rate was another hyperparameter optimized. Although not part of the architecture itself, the learning rate is a crucial hyperparameter for any learning algorithm that was also optimized.
3.2 Convolutional neural networks
Although designed for and used mainly with images, convolutional neural networks (CNNs) are suitable for any kind of grid-like data, especially when there is a natural fixed notion of distance between neighboring data points, such as the case for time series [GBC16].
Now, a ‘chunk’ of our dataset – a subtable of fixed length (in time) can be passed as a single input to our model. Evaluating the history from the last, say, $15$ days explicitly, the model is making a prediction for the following day as before.
Both one- and two-dimensional CNNs can be used for our problem. One-dimensional (1D) CNNs only perform the convolution operation over the time dimension, treating separate features as parallel input channels, just like the red, green and blue channels are considered when working with images. Two-dimensional (2D) CNNs take each chunk as a single-channel ‘image’, performing the convolution operation over both the time dimension and across the various features. The latter seems counterintuitive as there is no fixed distance between the features, but [HH19] found success with this technique. After experimenting with 2D CNNs on our data, we found them a lot more computationally expensive than others, while often performing worse. For this reason 2D CNNs were later dropped from our consideration.
The length in time of the chunks passed as individual examples (‘chunk length’) was a crucial hyperparameter to optimize for all of the CNNs. The number of convolutional layers (including pooling and activation functions) was fixed at $3$ for all CNN architectures to reduce the size of the search spaces, but the kernel sizes were optimized. The number of convolution filters per layer was optimized too, but it works differently for different implementations, as discussed below. Dropout layers were used once again for regularization and the dropout rate was optimized too, together with the learning rate.
3.2.1 Depthwise 1D CNNs
Among 1D CNNs, there are two distinct ways to perform the convolution operation. In depthwise convolution, each channel (input feature) is passed separately through its own convolutional and pooling layers. Each of these then gives a fixed number (the number of separate convolution filters) of output channels per input channel, interpreted as a summary of what happened to this feature in the time span considered. The convolution and pooling operations will shrink the length (in time) of each output channel, possibly to $1$ (otherwise flattening is applied), and all of these are then passed to a single dense layer that makes a prediction based on these summaries.
Although easy to interpret, this method had the problem of blowing up the dimensionality of the problem, as each of the features gets its own set of convolution parameters. This made computation infeasible and more work is needed to be carried out for this architecture.
3.2.2 Ordinary 1D CNNs
In basic (we call them ‘ordinary’) 1D convolution, all input channels are mixed already in the first (and subsequent) convolution layers. The output of convolution is flattened and passed to a single dense layer for making the final prediction.
3.3 Recurrent neural networks
Recurrent neural networks (RNNs) were designed for and are very successful at treating sequential data [GBC16] and are thus the natural choice for our problem. Both simple RNNs and more sophisticated long short-term memory (LSTM) RNNs were implemented and optimized separately, but they have the same search spaces of hyperparameters. In both cases, stacked RNNs were used, where the output of one RNN is taken as input to the next one. The number of such stacked layers was a hyperparameter to optimize. Each RNN layer also has its number of hidden units, which was also optimized. The chunk length, here naturally interpreted as the length of the input sequence, was once again optimized, together with dropout and learning rates as before.
LSTMs provide many advantages over basic RNNs [GBC16] without too much additional computational cost, especially in the case of our small datasets. For this reason after initial trials, we proceeded with just LSTMs for the final tests.
3.4 Temporal fusion transformer
Forecasting across multiple time horizons often involves a intricate amalgamation of inputs, encompassing static (i.e., time-invariant) covariates, known future inputs, and additional exogenous time series observed solely in the past. The challenge lies in handling this complexity without prior knowledge of how these inputs interact with the target variable. While various deep learning approaches have been proposed, they often manifest as ‘black-box’ models, lacking transparency regarding their utilization of the diverse input types encountered in practical scenarios.
In a recent work [LALP21], the Temporal Fusion Transformer (TFT) was introduced as an innovative attention-based architecture addressing this issue. TFT not only achieves high-performance multi-horizon forecasting but also provides interpretable insights into temporal dynamics. In order to capture temporal relationships at different scales, TFT incorporates recurrent layers for local processing and interpretable self-attention layers for modelling long-term dependencies. Specialized components within TFT are employed to select relevant features, and a series of gating layers effectively suppress unnecessary components, resulting in high performance across a broad spectrum of scenarios.
[LALP21] demonstrated significant performance enhancements over existing benchmarks across various real-world datasets. The study also highlighted three practical use cases illustrating the interpretability of TFT, showcasing its efficacy in shedding light on the decision-making process. The PyTorch Forecasting package [PyT23] has on open-source implementation of the model. It comes with its own optimizer, selecting the optimal number of attention heads, the network’s hidden size, and learning and dropout rates.
Although very promising and successful for other datasets, we found performance of TFT rather poor on our data: the models reduce to a trivial binary predictor, always predicting $1$ or 00 no matter the input test data. We believe this was largely due to the lack of data. As previously mentioned, novel complex architectures like TFT have tens if not hundreds of thousands of parameters which need a lot of data to train well. It would be interesting to revisit TFT and other models derived from it that have recently appeared in the literature on other, bigger datasets.
4 Challenges and performance estimation
There are many evaluation metrics applicable to the binary classification problem. Although common for other classification tasks, accuracy ($ACC$) is not a good metric as a trivial predictor that always gives the same output no matter the input can have high accuracy in the case of an unbalanced dataset. Balanced accuracy ($bACC$), the average between the true negative and true positive rates, mitigates this problem. Another popular alternative is the $F_{1}$ score, the harmonic mean of precision and recall. The area under the receiver operating characteristic (ROC) curve, or $AUC$ for short is another good choice. It is hard to say which one is better, and a good strategy could mix a combination of these.
Predicting financial markets is extremely challenging and this is reflected in our test metrics always being very close to $0.5$, i.e. the predictors being very close to random. This would be regarded as bad performance in many areas, but for financial datasets it is quite the norm: many companies make successful investments using algorithms (whether traditional or ML-based) being only marginally better than a random predictor [HH19]. Further, neural networks in particular in this context suffer from comparatively high variance depending on the random seed chosen during training due to their stochastic nature [HH19]. Even identical neural network architectures, trained in exactly the same way but with different random seeds can have vastly different performance on the test dataset, e.g. $0.56$ on a ‘good’ random seed and $0.48$ (i.e. worse than random) on a ‘bad’ one. This is a problem for NAS as what seemed like a good (or bad) neural network configuration might have just been a lucky (or unlucky respectively) seed. To mitigate this problem, we trained and tested each architecture configuration chosen $15$ times with different random seeds, averaging the test metrics and using the averages as our guide. At the end of optimization, the best performing configuration was trained and tested $50$ times (again, with different random seeds every time) to further eliminate random seed variation.
5 Search strategies
We started with $7$ different broad architecture types: FFNN, depthwise 1D CNN, ordinary 1D CNN, 2D CNN, simple RNN, LSTM, and TFT, but for final optimization studies only FFNN, ordinary 1D CNN, and LSTM were used, as discussed in Section 3. Each of them have their own search space of hyperparameters determining the exact network architecture.
Classical approaches for hyperparameter tuning in machine learning, such as grid search and random search, have been widely used due to their simplicity [BB12]. However, these methods can be computationally expensive and inefficient, especially when dealing with high-dimensional search spaces or expensive objective functions [BBBK11]. The survey on NAS [${}^{+}$23] favourably featured three other popular search strategies: Bayesian optimization, reinforcement learning, and the hyperband method which are the three methods we explored extensively.
Fundamentally, all of these suggest a specific configuration of hyperparameters of a neural network. The network is then trained and tested (in our case repeatedly so, see Section 4), the results saved, and the algorithm suggests the next configuration to try. This continues for a set number of trials (in our studies, we set it at $300$) and in the end the best performing configuration is selected.
5.1 Bayesian optimization
Bayesian optimization has emerged as a popular alternative to grid and random search for hyperparameter tuning, as it efficiently explores the search space and intelligently guides the optimization process [SLA12]. One widely used technique in Bayesian optimization is based on Gaussian Processes (GP), which fit the objective function using Gaussian processes as probabilistic models and leverages acquisition functions to balance exploration and exploitation [RW06]. GPs provide a flexible, non-parametric tool for modeling complex functions, allowing for uncertainty quantification and adaptation to new data [${}^{+}$16].
However, GP-based methods have some limitations, such as the inability to handle categorical features or dependent parameters without ad hoc modifications [BCdF10]. This has led to the development of alternative methods, such as the Tree-structured Parzen Estimator (TPE), which has gained attention for its scalability and effectiveness [BBBK11]. TPE models the search space using hierarchical Parzen estimators, which adaptively partition the space to model complex, high-dimensional functions [BBBK11].
TPE is particularly suitable for addressing the shortcomings of GP-based methods, as it can naturally handle categorical features and dependent parameters without requiring extensive adaptations. Performance comparisons between TPE, GP-based methods, and classical search techniques have shown that TPE often yields improved results across various tasks and datasets [${}^{+}$13].
A notable implementation of TPE is available in the Optuna package, a versatile optimization library that has demonstrated its utility in various machine learning tasks [${}^{+}$19].
5.2 Reinforcement learning approach
The reinforcement learning method consists of treating the neural architecture search problem as a reinforcement learning problem. We construct an iterative approach where the parameters (and structure) of the neural architecture are initially chosen randomly, just as in the case of Bayesian optimization. After the neural network is constructed, it is trained for a number of periods using a train and validation time-series dataset, sometimes even a number of times in order to correctly assess its performance in predicting the time series, and exclude issues relating to the choice of initial random seed. At the end of the run, a reward function is calculated, which represents the incremental increase in performance. In this way choices of parameters which produce “good” neural networks are rewarded, whereas those which produce badly-performing neural networks are penalized. We let the system evolve based for a number of iterations (reinforcement learning trials) and at the end we obtain the best overall neural network configuration.
The specific method we used introduced in [ZL17] uses an RNN as a controller, trained with REINFORCE. An open-source implementation of the controller RNN to use with image data is available at [Maj18], which we adpated to work with our time series data.
The reinforcement learning approach has several advantages compared to the Bayesian optimization: it is able to capture all the dynamics and interdependencies between the parameters, something that is difficult with classic bayesian optimization without testing every possible combination of parameters. Moreover we can more easily use customized reward functions (for example, in one case the reward function used was the average of the AUC and F1 scores) which can better capture the performance of the neural network in the presence of imbalanced datasets.
5.3 Hyperband
Hyperband is a method introduced in 2018 by Li et al., bringing in a clever tactic known as ‘bandit’ to make the hyperparameter optimization process very effective [${}^{+}$18]. It explores various combinations, paying more attention to the promising ones and discarding the under-performers.
What makes Hyperband notable are its efficiency, scalability, and the ability to use multiple GPUs simultaneously for faster optimization. It is compatible with popular machine learning libraries like Scikit-learn and Keras, embedded in the KerasTuner framework [${}^{+}$19].
Hyperband is designed for the efficient exploration and optimization of hyperparameter configurations. It selectively targets promising models in the search space, optimizing computational resources and time. The algorithm employs successive halving, a subprocess that involves eliminating less promising models and further training the survivors.
In successive halving, you provide a list of models ($models$), the total epochs needed for full training ($max\_epochs$), the initial training epochs for all models ($start\_epochs$), and a factor determining the speed of reducing models and increasing training epochs for the survivors ($factor$). This process iterates until all models are fully trained, resulting in the identification of the best model and its validation loss.
To achieve a balance between exploration and optimization, Hyperband executes successive halving multiple times, labeling each iteration as a bracket. The number of brackets is determined by a parameter, $s\_max$, approximately equal to $log(max\_epochs,factor)$. In each bracket, the number of models decreases by a factor, and new models are generated. Hyperband ensures a robust exploration by randomly generating models for each bracket, preventing too few models in a bracket. The initial training epochs for each bracket strictly increase by a factor.
Practically implemented in KerasTuner as the Hyperband tuner, it is available for single or multiple GPUs for parallel tuning. Users can specify $factor$ and $max\_epochs$, along with the number of brackets using $hyperband\_iterations$. This allows control over the search time while efficiently exploring the hyperparameter space [SJH22].
6 Results
The data was split into (in historical order): first $70\%$ for training, the following $20\%$ for validation used in the evaluation during optimization runs, and the last $10\%$ for final testing on unseen data. When the optimization studies were completed, the architectures performing best on the validation dataset were selected and evaluated on the test set. This involved retraining and testing each network $50$ times, and average performance was compared. This was done for each search strategy, architecture type and dataset combination.
Our best results were obtained on the Germany dataset, with time-derived redundant features removed but no PCA applied. The LSTM with parameters selected by the hyperband method applied to the unseen test data achieved an $AUC$ score of $0.56$ on average over $50$ test runs starting from different random seeds (see Section 4), with $0.05$ standard deviation. The same architecture showed balanced accuracy score of $0.54\pm 0.04$. 1D CNNs were also able to achieve good performance on this dataset, with the best architecture provided by Bayesian optimization giving an $AUC$ score of $0.54\pm 0.05$ and a high $F1$ score of $0.6\pm 0.06$.
For the Japan dataset, the redundant time-derived features removed and PCA applied retaining most of the information in the dataset while further cutting the number of features approximately in half gave the best results. The best performing architecture was a 1D CNN coming from Bayesian optimization, achieving an $AUC$ score of $0.54\pm 0.03$ over $50$ test runs, with a high $F1$ score of $0.65\pm 0.02$.
The US dataset was most challenging for us. Although tuned 1d CNNs gave an AUC score of $0.6\pm 0.02$ on validation data (not used for training) in our repeated testing, and an individual lucky random seed gave an $AUC$ of $0.58$ on test data, on average no architecture could achieve $AUC$ over $0.5$ on the test dataset. We believe something very significant happened in the US market in the last year or so of our data that changed market dynamics.
6.1 Architectures and search strategies compared
Our results show LSTMs and 1D CNNs outperforming simple feedforward networks for nearly all search strategies and datasets, as expected, but were neck in neck between each other.
Figure 1: Best performing architectures selected by each search strategy on each dataset. Every point represents average $AUC$ score and standard deviation on test data after retraining the selected architecture $50$ times. Type of the neural network chosen by the search strategy is displayed above each point, search strategies are color-coded.
For search strategies, overall the hyperband method and bayesian optimization showed better performance than the reinforcement learning based approach, but all three were very close to each other, well within standard deviation as shown in Figure 1. Optimization time was very similar between the three methods, mostly determined by the architecture type, total number of trials, repetitions per trial (to minimize variation due to different random seeds), and the number of epochs to train each neural network. These were set manually by us to 300, 15, and 80 respectively, for a total optimization time of about 12 hours per architecture type using a single Nvidia Quadro RTX 4000 GPU, with feedforward networks training faster than CNNs and LSTMs due to their lower complexity.
Practically, the reinforcement learning approach required the most work to adapt from the open-source implementation of the method in [ZL17], and has many parameters of its own to be optimized for best performance. In contrast, both the hyperband implementation in KerasTuner and the Bayesian optimization from Optuna are easy to use and well-supported. The Optuna implementation is flexible in that nearly the same code can be used to optimize hyperparameters of a different kind of solver other than neural networks, such as decision trees-based algorithms popular for time series prediction. Meanwhile, KerasTuner implementation of hyperband has advantage of easy parallelization over several GPUs (although we didn’t have the hardware to make use of it).
7 Discussion
Overall, we didn’t observe significant convergence over time in the hyperparameter optimization. An example history of optimization plot is shown in Figure 2. This was a general problem present for all neural architecture types, optimization strategies and datasets.
Figure 2: Bayesian optimization history for the ordinary 1D CNN architecture on the US dataset. Each point represents average $AUC$ score on the validation dataset over $15$ runs for the same network configuration.
We believe all search strategies were working as they should, but the main problem of financial time series prediction on our datasets was just too difficult to decisively improve upon through neural architecture search alone. Many steps were taken to improve overall performance. In the beginning the networks were heavily overfitting, where we struggled to get any metrics above $0.5$ even on the validation data. The feature reduction measures such as removing time-derived features and PCA discussed in Section 2 helped mitigate this. Unfortunately, this still did not always translate to good performance on the test dataset, even though we heavily used dropout layers for regularization. We also tried converting the main binary clasification problem to regression, but this did not result in substantial improvement.
As a consequence, simply taking the final suggested architecture by either search strategy didn’t give the best results. The best performing ones were also not always the best choice due to the random seed variance discussed in Section 4. Even though each network configuration was trained and tested $15$ times during the optimization process, training and testing the best architectures more (e.g., 50) times sometimes showed worse performance, indicating that the issue of random seed variation was not fully resolved in our approach.
Nevertheless, all search strategies were still successful in highlighting best choices for network parameters, at least for the validation set, which is most evident from slice plots as in Figure 3. For bayesian optimization, the best parameters for each architecture were read off slice plots like this. However we note that following this approach somewhat negates the main advantages in going for more advanced search strategies as compared to simple random search.
Figure 3: Slice plot for bayesian optimization of LSTMs on the US dataset. Each point represents average $AUC$ score on the validation dataset over $15$ runs for the same network configuration. It is clear that setting chunk_length (i.e. the length in time of a sequence passed to the LSTM for individual prediction) to $10$ gives the best results, while other parameters are less impactful.
To combat the big variance between networks starting from different random seeds (see Section 2 and Figure 1), we tried filtering the outputs, disregarding those with probabilities close to $0.5$ (interpreted as those where the network ‘was not sure’), a standard approach [GBC16]. Although this improved overall performance a little on average, it instead greatly increased the variance. For a fixed architecture, selecting networks trained from random seeds giving good performance on validation set did not translate to better performance on the test set. Our best suggestion to combat the variance is to use an ensemble model built from many incarnations of a chosen network architecture.
8 Future work
Performing the same studies on more and publicly available time-series datasets, especially non-financial data would shed more light whether our findings were specific to our data or a more general feature.
In addition, cell-based and hierarchical search spaces [${}^{+}$23] could be explored in the future in the context of time-series forecasting. These are based on searching over structural blocks within a network piece by piece instead of navigating through a grid-like space as in the standard approach we took. Finally, one-shot NAS techniques such as supernet-based methods (differentiable and not) and hypernetworks [${}^{+}$23] could be explored. With one-shot techniques, one trains a single (massive) super/hyper-network, which can be used to subsample and evaluate many smaller networks to find the optimal one, without re-training. This is opposed to the standard approach of training and testing many architecture configurations that we took.